# Pluriclosed metrics on compact semisimple Lie groups

### v1

@@ Abstract.
Given a compact semisimple Lie group $G$ and a maximal torus $T\subset G$, we give an explicit description of all left and $\operatorname{Ad}(T)$-invariant pluriclosed Hermitian structures on $G$ in terms of the corresponding root system. They depend on $12d+$ 1 parameters in the irreducible case, where $\dim{T}=2d$.
@@ 1. Introduction
A Hermitian metric $g$ on a complex manifold $(M,J)$ is said to be pluriclosed (or SKT, short for strong Kähler with torsion) if the Kähler form $\omega=g(J\cdot,\cdot)$ satisfies that $dd^{c}\omega=0$ (we refer to the recent surveys [F, FG2] for further information). Being $dd^{c}\omega$ a $4$-form, the condition turns out to be very strong in high dimensions, specially in the homogeneous case, making the search for examples a hard task (see [AL, AN, FP, FS] for structure and classification results on solvmanifolds). The primal compact examples are bi-invariant metrics on any even dimensional compact semisimple Lie group $G$ endowed with any left-invariant complex structure. As known, these are precisely Bismut flat manifolds with finite fundamental group (see [WYZ]). It was proved in [FGV] that a compact homogeneous complex manifold with finite fundamental group (i.e., a C-space) admits an invariant pluriclosed metric only if it is a Lie group $G$ (modulo a flag manifold, i.e., a Kähler homogeneous manifold).
Any left-invariant complex structure $J$ on a compact semisimple Lie group $G$ is determined by the choices of a maximal torus $T\subset G$ and a set of positive roots $\Delta^{+}$ (see [S, W, Pi]), they are often called Samelson complex structures. Indeed, $J$ leaves invariant the $\operatorname{B}_{\mathfrak{g}}$-orthogonal decomposition $\mathfrak{g}=\mathfrak{t}\oplus\mathfrak{q}$ of the Lie algebra $\mathfrak{g}$ of $G$, where $\operatorname{B}_{\mathfrak{g}}$ is the Killing form of $\mathfrak{g}$, $J_{\mathfrak{t}}$ is any linear map such that $J_{\mathfrak{t}}^{2}=-I$ and on the complexification $\mathfrak{q}^{c}=\bigoplus\limits_{\alpha\in\Delta}{\mathbb{C}}E_{\alpha}$, the complex linear map determined by $J_{\mathfrak{q}}$ on $\mathfrak{q}^{c}$ is given by $J_{\mathfrak{q}}E_{\alpha}=\pm\mathrm{i}E_{\alpha}$ for all $\alpha\in\pm\Delta^{+}$. The space of all left-invariant complex structures on $G$ therefore depends on $2d^{2}$ parameters up to biholomorphism if $\dim{T}=2d$. Any left and $\operatorname{Ad}(T)$-invariant metric which is compatible with $J=(J_{\mathfrak{t}},J_{\mathfrak{q}})$ is of the form $g=g_{\mathfrak{t}}+g_{\mathfrak{q}}$, where $g_{\mathfrak{t}}$ is any inner product on $\mathfrak{t}$ compatible with $J_{\mathfrak{t}}$ and $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta^{+}}$ is given by
@@
    If $(G,J)$ admits a pluriclosed left-invariant metric, then it also admits a left and $\operatorname{Ad}(T)$-invariant pluriclosed metric $g$ for the maximal torus $T$ of $G$ attached to $J$.
@@
  4. (d)
    The only left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics which are also CYT (Calabi-Yau with torsion, i.e., the Bismut Ricci form vanishes) are bi-invariant metrics.
As far as we know, beyond $\mathrm{SU}(2)\times\mathrm{SU}(2)$, $\mathrm{SU}(3)$ and $\mathrm{SO}(9)$, the only other known examples of pluriclosed left-invariant metrics (non bi-invariant) on compact semisimple Lie groups were exhibited in [P] for $G=\mathrm{G}_{2}$. The pluriclosed flow in this setting was studied in [B], obtaining as an application the non-existence of pluriclosed and CYT metrics beyond the bi-invariant (or Bismut flat) ones.
In this paper, after giving a formula for $dd^{c}\omega$ in terms of roots for any left and $\operatorname{Ad}(T)$-invariant Hermitian structure on $G$ (see Lemma 3.1), we provide an alternative algebraic proof for the result (b) above (see Proposition 3.4) and then we prove the following complete classification (see Theorem 3.5 for a more detailed statement). We fix a left-invariant complex structure $J$ attached to a maximal torus $T$ and consider the decomposition $\mathfrak{t}=\mathfrak{t}_{1}\oplus\dots\oplus\mathfrak{t}_{s}$, where the $\mathfrak{t}_{i}$ are the corresponding maximal tori of the simple factors of $G$.
@@
Given a compact semisimple Lie group $G$ of rank $2d$ endowed with $J$, any left and $\operatorname{Ad}(T)$-invariant pluriclosed metric is given as follows: $g_{\mathfrak{t}}(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$ and on each simple factor, up to scaling, $g_{\mathfrak{t}}=-\operatorname{B}_{\mathfrak{g}}|_{\mathfrak{t}}$ and $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta^{+}}$, where
$x_{\alpha}:=\sum_{i=1}^{2d}k_{i}x_{i}-\left(\sum_{i=1}^{2d}k_{i}-1\right),\qquad\forall\alpha=\sum_{i=1}^{2d}k_{i}\alpha_{i}\in\Delta^{+},\quad k_{i}\in{\mathbb{Z}}_{\geq 0},$
$\Pi=\{\alpha_{1},\dots,\alpha_{2d}\}\subset\Delta^{+}$ is the set of simple roots and $x_{i}:=x_{\alpha_{i}}$ for $i=1,\dots,2d$ (equivalently, $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}-1$ for all $\alpha,\beta,\alpha+\beta\in\Delta^{+}$).
Each of these metrics is compatible with a $d(d-1)$-parametric space of left-invariant complex structures on $G$. According to the theorem, for a fixed $J$, the space of all left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics compatible with $J$ depends on $r(2d+1)$ parameters and contains an $r$-parametric space of bi-invariant metrics, where $r$ is the number of irreducible factors of $(G,J)$ as a complex manifold. In particular, any irreducible $J$ admits a unique compatible bi-invariant metric up to scaling. Note that $J$ can be irreducible on a non-simple $G$ (see Example 2.1).
@@
$x_{\alpha}:=\sum\limits_{i=1}^{2d}k_{i}x_{i},\qquad\forall\alpha=\sum\limits_{i=1}^{2d}k_{i}\alpha_{i}\in\Delta^{+}.$
@@
Potential applications of Theorem 1.1 include an alternative algebraic proof for result (d) above and the study of the pluriclosed flow as an ODE system for $(x_{1}(t),\dots,x_{2d}(t))$, which will be the object of further research.
Acknowledgements. We thank Beatrice Brienza for fruitful conversations.
@@ 2. Hermitian geometry of Lie groups
Let $G$ be a connected and compact semisimple Lie group of even dimension. The rank of $G$ is therefore also even, say $\operatorname{rk}(G)=2d$. We refer to §4 for the definitions and notations concerning the root system of $G$.
@@
In particular, $x_{\alpha}>0$ and we set $x_{-\alpha}:=x_{\alpha}$ for any $\alpha\in\Delta^{+}$. The corresponding symmetric ${\mathbb{C}}$-bilinear form on $\mathfrak{g}^{c}$, also denoted by $g$, satisfies that
@@
We note that $(G,J)$ is irreducible as a complex manifold if and only if $J_{\mathfrak{t}}$ does not leave any subspace $\mathfrak{t}_{i_{1}}\oplus\dots\oplus\mathfrak{t}_{i_{k}}$ with $0<k<s$ invariant, where $\mathfrak{t}_{i}$ is the maximal torus of $\mathfrak{g}_{i}$. In that case, there exists a unique bi-invariant metric compatible with $J$ (up to scaling), which is not necessarily the Killing metric as the following example shows.
@@
In particular, $d\omega$ is always nonzero, in accordance with the well-known fact that $(J,g)$ is never Kähler. It also follows that the Hermitian structure $(J_{\mathfrak{q}},g_{\mathfrak{q}})$ on the flag manifold $G/T$ is Kähler if and only if $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}$ for all $\alpha,\beta,\alpha+\beta\in\Delta^{+}$.
@@
3. Pluriclosed structures
@@
We are now ready to prove the main result of this paper, a complete classification of left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics on $G$. If $G=G_{1}\times\dots\times G_{s}$, where $G_{i}$ is simple for all $i$, then for each $1\leq i\leq s$ we consider the decomposition $\mathfrak{g}_{i}=\mathfrak{t}_{i}\oplus\mathfrak{q}_{i}$ such that $\mathfrak{t}=\mathfrak{t}_{1}\oplus\dots\oplus\mathfrak{t}_{s}$ and $\mathfrak{q}=\mathfrak{q}_{1}\oplus\dots\oplus\mathfrak{q}_{s}$, so each $\mathfrak{t}_{i}$ is a maximal abelian subalgebra of $\mathfrak{g}_{i}$ and define a root system $\Delta_{i}$ with $\Delta=\Delta_{1}\sqcup\dots\sqcup\Delta_{s}$.
@@
    $g_{\mathfrak{t}}(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$ and for each $1\leq i\leq s$, there exists $z_{i}>0$ such that $g_{\mathfrak{t}}|_{\mathfrak{t}_{i}}=z_{i}g_{\operatorname{B}}|_{\mathfrak{t}_{i}}$ and $g_{\mathfrak{q}}|_{\mathfrak{q}_{i}}=(z_{i}x_{\alpha})_{\alpha\in\Delta^{+}_{i}}$, where the $x_{\alpha}$ are defined as follows: if $\Pi=\{\alpha_{1},\dots,\alpha_{2d_{i}}\}$ are the simple roots of $\mathfrak{g}_{i}$ and we set $x_{j}:=x_{\alpha_{j}}$ for $j=1,\dots,2d_{i}$, then
    $x_{\alpha}:=\sum_{j=1}^{2d_{i}}k_{j}x_{j}-\left(\sum_{j=1}^{2d_{i}}k_{j}-1\right),\qquad\forall\alpha=\sum_{j=1}^{2d_{i}}k_{j}\alpha_{j}\in\Delta^{+}_{i},\quad k_{j}\in{\mathbb{Z}}_{\geq 0}.$
@@
The natural number $h(\alpha):=\sum\limits_{i=1}^{n}k_{i}$ is called the height of the root $\alpha$. In order to obtain a metric it is of course necessary that $x_{\alpha}>0$ for all $\alpha\in\Delta^{+}$, which for instance holds if $x_{i}\geq 1-\tfrac{1}{h(\alpha_{max})}$ for all $i=1,\dots,n$, where $\alpha_{max}$ is the maximal root.

### v2

@@ Abstract.
Given a compact semisimple Lie group $G$ and a maximal torus $T\subset G$, we give an explicit description of all left and $\operatorname{Ad}(T)$-invariant pluriclosed Hermitian structures on $G$ in terms of the corresponding root system. They depend on $12d+$ 1 parameters in the irreducible case, where $\dim{T}=2d$. As applications, we obtain that the only left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics which are also CYT are bi-invariant metrics (i.e., Bismut flat) and study the pluriclosed flow as a neat ODE system.
@@ 1. Introduction
A Hermitian metric $g$ on a complex manifold $(M,J)$ is said to be pluriclosed (or SKT, short for strong Kähler with torsion) if the Kähler form $\omega=g(J\cdot,\cdot)$ satisfies that $dd^{c}\omega=0$ (we refer to the recent surveys [F, FG2] for further information). Being $dd^{c}\omega$ a $4$-form, the condition turns out to be very strong in high dimensions, specially in the homogeneous case (see [AL, AN, FP, FS] for structure and classification results on solvmanifolds). The primal compact examples are bi-invariant metrics on any even dimensional compact semisimple Lie group $G$ endowed with any left-invariant complex structure. As known, these are precisely Bismut flat manifolds with finite fundamental group (see [WYZ]). It was proved in [FGV] that a compact homogeneous complex manifold with finite fundamental group (i.e., a C-space) admits an invariant pluriclosed metric only if it is a Lie group (modulo a flag manifold, i.e., a Kähler homogeneous manifold).
Any left-invariant complex structure $J$ on a compact semisimple Lie group $G$ is determined by the choices of a maximal torus $T\subset G$ and of a set of positive roots $\Delta^{+}$ (see [S, W, Pi]), they are often called Samelson complex structures. More precisely, $J$ leaves invariant the $\operatorname{B}_{\mathfrak{g}}$-orthogonal decomposition $\mathfrak{g}=\mathfrak{t}\oplus\mathfrak{q}$ of the Lie algebra $\mathfrak{g}$ of $G$, where $\operatorname{B}_{\mathfrak{g}}$ is the Killing form of $\mathfrak{g}$, $\mathfrak{t}$ is the Lie algebra of $T$, $J_{\mathfrak{t}}$ is any linear map such that $J_{\mathfrak{t}}^{2}=-I$ and on the complexification $\mathfrak{q}^{c}=\bigoplus\limits_{\alpha\in\Delta}{\mathbb{C}}E_{\alpha}$, the complex linear map determined by $J_{\mathfrak{q}}$ on $\mathfrak{q}^{c}$ is given by $J_{\mathfrak{q}}E_{\alpha}=\pm\mathrm{i}E_{\alpha}$ for all $\alpha\in\pm\Delta^{+}$. The space of all left-invariant complex structures on $G$ therefore depends on $2d^{2}$ parameters up to biholomorphism if $\dim{T}=2d$. Any left and $\operatorname{Ad}(T)$-invariant metric which is compatible with a fixed $J=(J_{\mathfrak{t}},J_{\mathfrak{q}})$ is of the form $g=g_{\mathfrak{t}}+g_{\mathfrak{q}}$, where $g_{\mathfrak{t}}$ is any inner product on $\mathfrak{t}$ compatible with $J_{\mathfrak{t}}$ and $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta^{+}}$ is given by
@@
    The only left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics which are also CYT (Calabi-Yau with torsion, i.e., the Bismut Ricci form $\rho^{B}$ vanishes, see §4) are bi-invariant metrics.
As far as we know, beyond $\mathrm{SU}(2)\times\mathrm{SU}(2)$, $\mathrm{SU}(3)$ and $\mathrm{SO}(9)$, the only other known examples of pluriclosed left-invariant metrics (non bi-invariant) on compact semisimple Lie groups were exhibited in [Ph] for $G=\mathrm{G}_{2}$.
In this paper, after giving a formula for $dd^{c}\omega$ in terms of roots for any left and $\operatorname{Ad}(T)$-invariant Hermitian structure on $G$ (see Lemma 3.1), we provide an alternative algebraic proof for the result stated in (a) above (see Proposition 3.6) and then we prove the following complete classification (see Theorem 3.7 for a more detailed statement). We fix a left-invariant complex structure $J$ attached to a maximal torus $T$ and consider the decomposition $\mathfrak{t}=\mathfrak{t}_{1}\oplus\dots\oplus\mathfrak{t}_{s}$, where the $\mathfrak{t}_{i}$ are the corresponding maximal tori of the simple factors of $G$.
@@
On a compact semisimple Lie group $G$ endowed with a Samelson complex structure $J$, any left and $\operatorname{Ad}(T)$-invariant pluriclosed metric is given as follows: $g_{\mathfrak{t}}(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$ and on each simple factor, up to scaling, $g_{\mathfrak{t}}=-\operatorname{B}_{\mathfrak{g}}|_{\mathfrak{t}}$ and $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta^{+}}$, where
$x_{\alpha}:=\sum_{i=1}^{n}k^{\alpha}_{i}x_{i}-\left(\sum_{i=1}^{n}k^{\alpha}_{i}-1\right),\qquad\forall\alpha=\sum_{i=1}^{n}k^{\alpha}_{i}\alpha_{i}\in\Delta^{+},\quad k^{\alpha}_{i}\in{\mathbb{Z}}_{\geq 0},$
$\Pi=\{\alpha_{1},\dots,\alpha_{n}\}\subset\Delta^{+}$ is the set of simple roots and $x_{i}:=x_{\alpha_{i}}$ for $i=1,\dots,n$ (equivalently, $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}-1$ for all $\alpha,\beta,\alpha+\beta\in\Delta^{+}$).
If $\operatorname{rk}(G)=2d$, then each of these metrics is compatible with a $d(d-1)$-parametric space of left-invariant complex structures on $G$. According to the theorem, for a fixed $J$, the space of all left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics compatible with $J$ depends on $r(2d+1)$ parameters and contains an $r$-parametric space of bi-invariant metrics, where $r$ is the number of irreducible factors of $(G,J)$ as a complex manifold. In particular, any irreducible $J$ admits a unique compatible bi-invariant metric up to scaling. Note that $J$ can be irreducible on a non-simple $G$ (see Example 2.1).
@@
$x_{\alpha}:=\sum\limits_{i=1}^{n}k^{\alpha}_{i}x_{i},\qquad\forall\alpha=\sum\limits_{i=1}^{n}k^{\alpha}_{i}\alpha_{i}\in\Delta^{+}.$
@@
As a first application of the explicit description of pluriclosed metrics given in Theorem 1.1, we provide in §4 an alternative proof for the result stated in (c) above using the convexity properties of the functional
(1) $F(g)=F(x_{1},\dots,x_{n}):=\sum_{\alpha\in\Delta^{+}}x_{\alpha}-\log{x_{\alpha}}.$
Secondly, we apply Theorem 1.1 to study the pluriclosed flow $\tfrac{\partial}{\partial t}\omega=-(\rho^{B})^{1,1}$ for a one-parameter family $\omega(t)=g(t)(J\cdot,\cdot)$ as an ODE system for
$X(t):=(x_{1}(t),\dots,x_{n}(t))\in{\mathbb{R}}^{n}.$
The pluriclosed flow was introduced in [ST] and in the homogeneous case, it has been studied mostly on solvmanifolds (see [AL, FV] and the references therein). On a compact semisimple Lie group $G$, using a general convergence result from [GJS], Barbaro proved in [B] the convergence to a Bismut flat metric starting at any pluriclosed metric (not necessarily left-invariant) by computing the $(1,1)$-Aeppli cohomology of $G$, obtaining as an application the non-existence of pluriclosed and CYT metrics on $G$ other than the bi-invariant ones.
We show in §5 that the metrics $g(t)$ attached to $X(t)$ pluriclosed flow according to the following ODE system:
(2) $x_{j}^{\prime}=\sum_{\alpha\in\Delta^{+},1\leq i\leq n}\left(1-\tfrac{1}{x_{\alpha}}\right)k^{\alpha}_{i}\langle\alpha_{i},\alpha_{j}\rangle,\qquad\forall j=1,\dots,n,$
or equivalently,
(3) $X^{\prime}(t)=-Q\operatorname{grad}(F)_{X(t)},\qquad\mbox{where}\quad Q:=[\langle\alpha_{i},\alpha_{j}\rangle],$
and $F$ is the functional defined in (1). This converts $F$ in a Lyapunov function for the dynamical system in the sense that $F(X(t))$ is strictly decreasing along any solution $X(t)$, which implies that any $X(t)$ converges to $(1,\dots,1)$ as $t\to\infty$. Any left and $\operatorname{Ad}(T)$-invariant pluriclosed solution $g(t)$ therefore converges, as $t\to\infty$, to a bi-invariant metric (see Theorem 5.1 for a more precise statement). We note that the ODE systems (2) and (3) allows us to actually see the pluriclosed in action among a large class of metrics. This may for instance be used to establish the evolution behavior of many geometric quantities along these pluriclosed flow solutions.
Acknowledgements. We thank Beatrice Brienza and Ramiro Lafuente for fruitful conversations and to Fabio Podestà for bringing the paper [AD] to our attention.
@@ 2. Hermitian geometry of Lie groups
Let $G$ be a connected and compact semisimple Lie group of even dimension. The rank of $G$ is therefore also even, say $\operatorname{rk}(G)=2d$. We refer to §6 for the definitions and notation concerning the root system of $G$.
@@
Equivalently,
$\left\{\tfrac{1}{\sqrt{x_{\alpha}}}e_{\alpha},\;\tfrac{1}{\sqrt{x_{\alpha}}}J_{\mathfrak{q}}e_{\alpha}:\alpha\in\Delta^{+}\right\}$
is a $g$-orthonormal basis of $\mathfrak{q}$. In particular, $x_{\alpha}>0$ and we set $x_{-\alpha}:=x_{\alpha}$ for any $\alpha\in\Delta^{+}$. The corresponding symmetric ${\mathbb{C}}$-bilinear form on $\mathfrak{g}^{c}$, also denoted by $g$, satisfies that
@@
We note that $(G,J)$ is irreducible as a complex manifold if and only if $J_{\mathfrak{t}}$ does not leave any subspace $\mathfrak{t}_{i_{1}}\oplus\dots\oplus\mathfrak{t}_{i_{k}}$ with $0<k<s$ invariant, where $\mathfrak{t}_{i}$ is the maximal torus of $\mathfrak{g}_{i}$ (cf. [B, Section 2]). In that case, there exists a unique bi-invariant metric compatible with $J$ (up to scaling), which is not necessarily the Killing metric as the following example shows.
@@
In particular, $d\omega$ is always nonzero, in accordance with the well-known fact that $(J,g)$ is never Kähler. It also follows that the Hermitian structure $(J_{\mathfrak{q}},g_{\mathfrak{q}})$ on the full flag manifold $G/T$ is Kähler if and only if $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}$ for all $\alpha,\beta,\alpha+\beta\in\Delta^{+}$.
@@
Left-invariant $1$-forms and closed $2$-forms on $G$ are all respectively given by
$\theta_{X}:=g(\cdot,X),\qquad\sigma_{X}:=\operatorname{B}_{\mathfrak{g}}([\cdot,\cdot],X),\quad X\in\mathfrak{g},$
and using that $\sigma_{PX}=d\theta_{X}$ for all $X$, where $g=g_{\operatorname{B}}(P\cdot,\cdot)$, one obtains the well-known fact that the first and second Betti numbers of $G$ both vanish (see e.g. [LW, Section 2]).
The following vectors in $\mathfrak{t}$ play a crucial role in the geometry of the Hermitian manifold $(G,J,g)$ (see e.g. [AP, G, K, P]):
(4) $Z_{J_{\mathfrak{q}}}:=\sum_{\alpha\in\Delta^{+}}\mathrm{i}H_{\alpha},\qquad Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}}:=\sum_{\alpha\in\Delta^{+}}\tfrac{1}{x_{\alpha}}\mathrm{i}H_{\alpha}.$
Note that $Z_{J_{\mathfrak{q}}}=Z_{J_{\mathfrak{q}},g_{\operatorname{B}}|_{\mathfrak{q}}}$.
Lemma 2.4.
For any left and $\operatorname{Ad}(T)$-invariant Hermitian structure $(J,g)$ on a compact semisimple Lie group $G$,
$d_{g}^{*}\omega=-\theta_{Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}}}\quad\mbox{and}\quad dd_{g}^{*}\omega=-\sigma_{P_{\mathfrak{t}}Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}}},\quad\mbox{where}\quad g_{\mathfrak{t}}=g_{\operatorname{B}}(P_{\mathfrak{t}}\cdot,\cdot).$
Remark 2.5.
In particular, $d_{g}^{*}\omega=-g(\cdot,Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}})$ and
$dd_{g}^{*}\omega=-\operatorname{B}_{\mathfrak{g}}([\cdot,\cdot],P_{\mathfrak{t}}Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}})=g([\cdot,\cdot],Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}}).$
Proof.
For any $X\in\mathfrak{g}$, we have that
$d_{g}^{*}\omega(X)$ $=g(d_{g}^{*}\omega,\theta_{X})=g(\omega,d\theta_{X})=-\sum_{\alpha\in\Delta_{\mathfrak{q}}^{+}}\tfrac{1}{x_{\alpha}^{2}}\omega(e_{\alpha},J_{\mathfrak{q}}e_{\alpha})\theta_{X}([e_{\alpha},J_{\mathfrak{q}}e_{\alpha}])$
$=-\sum_{\alpha\in\Delta_{\mathfrak{q}}^{+}}\tfrac{1}{x_{\alpha}}\theta_{X}(\mathrm{i}H_{\alpha})=-\theta_{X}(Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}})=-g(X,Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}}),$
concluding the proof. ∎
3. Pluriclosed metrics
@@
Remark 3.3.
These formulas are in accordance with those given in [AD] for full flag manifolds.
@@
Remark 3.5.
This characterization coincides with one obtained in the case of a non-compact simple Lie group of inner type in [GP, (4.2)].
@@
We are now ready to prove the main result of this paper, a complete classification of left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics on $G$. If $G=G_{1}\times\dots\times G_{s}$, where $G_{i}$ is simple for all $i$, then for each $1\leq i\leq s$ we consider the decomposition $\mathfrak{g}_{i}=\mathfrak{t}_{i}\oplus\mathfrak{q}_{i}$ such that $\mathfrak{t}=\mathfrak{t}_{1}\oplus\dots\oplus\mathfrak{t}_{s}$ and $\mathfrak{q}=\mathfrak{q}_{1}\oplus\dots\oplus\mathfrak{q}_{s}$, so each $\mathfrak{t}_{i}$ is a maximal abelian subalgebra of $\mathfrak{g}_{i}$ and defines a root system $\Delta_{i}$ with $\Delta=\Delta_{1}\sqcup\dots\sqcup\Delta_{s}$.
@@
    $g_{\mathfrak{t}}(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$ and for each $1\leq i\leq s$, there exists $z_{i}>0$ such that $g_{\mathfrak{t}}|_{\mathfrak{t}_{i}}=z_{i}g_{\operatorname{B}}|_{\mathfrak{t}_{i}}$ and $g_{\mathfrak{q}}|_{\mathfrak{q}_{i}}=(z_{i}x_{\alpha})_{\alpha\in\Delta^{+}_{i}}$, where the $x_{\alpha}$ are defined as follows: if $\Pi_{i}=\{\alpha_{1},\dots,\alpha_{n_{i}}\}$ are the simple roots of $\mathfrak{g}_{i}$ and we set $x_{j}:=x_{\alpha_{j}}$ for $j=1,\dots,n_{i}$, then
    $x_{\alpha}:=\sum_{j=1}^{n_{i}}k^{\alpha}_{j}x_{j}-\left(\sum_{j=1}^{n_{i}}k^{\alpha}_{j}-1\right),\qquad\forall\alpha=\sum_{j=1}^{n_{i}}k^{\alpha}_{j}\alpha_{j}\in\Delta^{+}_{i},\quad k^{\alpha}_{j}\in{\mathbb{Z}}_{\geq 0}.$
@@
The natural number $h(\alpha):=\sum\limits_{i=1}^{n}k^{\alpha}_{i}$ is called the height of the root $\alpha=\sum\limits_{i=1}^{n}k^{\alpha}_{i}\alpha_{i}$, where $\Pi=\{\alpha_{1},\dots,\alpha_{n}\}$. In order to obtain a metric it is of course necessary that $x_{\alpha}>0$ for all $\alpha\in\Delta^{+}$, which for instance holds if $x_{i}\geq 1-\tfrac{1}{h(\alpha_{max})}$ for all $i=1,\dots,n$, where $\alpha_{max}$ is the maximal root.